OpenStudy (anonymous):

Someone please help me understand? (Will give medal) Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

Hi Aubrey- Its JO

OpenStudy (anonymous):

I see you.. haha

OpenStudy (anonymous):

Do you still need help?

jhonyy9 (jhonyy9):

f(1)=2 f(2)=3 f(3)=f(3-1) +f(3-2) = f(2) +f(1) = 3+2 =5 f(4) = f(4-1) +f(4-2) = f(3) +f(2) = 5 +3 = 8 f(5) = f(5-1) +f(5-2) = f(4) +f(3) = 8 +5 = 13 hope so much you can understanding it sure easy

OpenStudy (anonymous):

nope I got it @twistnflip I figured it out and thank you so much! @johnweldon1993 I understood what you have to do but thank you so much for answering(: